# SolNet: Open-source deep learning models for photovoltaic power forecasting across the globe

### v1

@@ 2.4 Model evaluation
Performance is measured by way of the RMSE, in accordance with (yang2020verification). yang2020verification state that the RMSE is favored over the MAE because it penalizes large errors more and because it is more popular. The first reason makes sense in solar forecasting since it is important to correctly predict the timing and size of the daily peak in output. The second reason is essentially a ’network effect’, where the value of using something increases when more people use it. This makes sense, since we want results to comparable over studies, so using the same error metric as most people use makes sense. In all cases, a full year of target data is kept aside for final evaluation. 80% of the remaining data is used for training and 20% is used to validate the model performance and pick the best performing iteration of the model.
@@
Figure 2: The workflow of the experiments
@@ 3 Data
Two types of data are required for this research: (1) the target domain data which has been observed in physical sites of interest but is limited in quantity, and (2) the source domain data, which is available in ample amounts, but has actually been synthetically generated from a (similar) setup. First, we briefly discuss the target domain datasets, i.e. PV installations from the Netherlands, Australia and Belgium, with only around two years of data per site. Second, we discuss the synthetic source domain used for transfer learning. Finally, we discuss the input features used in the forecast model, as well as the necessary pre-processing steps.
@@ 3.1.1 Residential buildings in the Netherlands
Our main target domain data comes from different solar PV systems installed in several locations, both at the residential and commercial scale. The main bulk of the data comes from 40 individual household PV systems spread across the Netherlands. The data comes at a temporal resolution of 15 minutes (which we resample to an hourly resolution in accordance with the source domain). Even though we have more data available from Australia, the Dutch dataset is seen as our main corpus because we have site-specific meta-data as well, including the tilt, the azimuth, the peak power, and the approximate location of the PV systems (zip code and city). From this latter, we derive an approximate latitude and longitude. The peak power of the systems varies between 1.2 kWp and 3.1 kWp, with most systems around 2.5 kWp.
Data quantity varies by location: for all 40 locations, the full year of 2020 is available. Most locations also have data for the entire year of 2019 (although in some cases, this can be as few as three months only). As mentioned, the full year of 2020 is kept aside for evaluation of the final models. This fits well with the core premise of the research that there is often only a (very) limited amount of data available in the target domain before forecasts need to be made.
@@ 3.1.2 Residential buildings in Australia
We also utilise solar data from 300 Australian homes with rooftop solar systems spanning the period of 1 July 2011 to 30 June 2013. The solar systems have a mean size of 1.68 kWp and are all located in New South Wales (specifically in zip codes between 20 and 23). No meta-data is available for these locations other than the size of the system and the zip code.
@@ 3.1.3 Commercial buildings in Belgium
We also consider PV production data from two commercial buildings in Genk, Belgium. Data is available for the full year of 2019 for one location, and for nine months in the other location (the entire year of 2020 is available for testing, as before). In this case, the exact location of the systems is well-known, but the tilt and azimuth of these systems are not well-specified, because individual panels at the same site are placed in a variety of configurations. The systems have a peak power of 275 kWp and 125 kWp.
@@ 3.2 Source domain
For this research we use data generated by PVGIS from 2005 to 2019, mimicking the parameters (location, tilt, azimuth, peak power) of the previously discussed solar PV systems (in case these parameters are available) in order to maximize the similarity between source and target domain.
This data is subsequently used to train the source models we use for transfer learning. These models are then fine-tuned using the sparse observational target data. Even though PVGIS provides data for 2020, we do not use this year while creating source models as this can potentially cause data leakage. Data leakage refers to including knowledge about the future (or data which would realistically not be available) in a model. In this case, data leakage would occur if we were to use the year 2020 in the creation of our model and, simultaneously, evaluate the model on 2020. Similarly, for Australia we only use data up until 2011.
3.3 Feature engineering
By only looking at lagged load values and not including any external covariates, we keep the model as general as possible, requiring no additional data from a potential user of the source models. However, we can use the fact that we are dealing with time series data to include additional cyclical variables. Given the diurnal and annual patterns of solar power generation, it makes sense to add variables that capture these patterns. This is done by doing a cyclical encoding of the hours and months, by creating a sine and cosine feature for each of them. Featurising the pattern in this manner has several advantages. First, the cyclicality between the periods is respected, which makes it preferable to numerical encoding. Second, this method only creates 2 additional variables for the pattern, which makes it preferable to one-hot encoding.
@@ 3.4 Data pre-processing
In order to properly train and evaluate the results of our models and avoid leakage, the data has to be split into several sets. The source data is split in a train and test set, while the target data is split in a train set, a validation set, and a separate test set. We do this three-way split to mimic a realistic scenario where a proprietor of the system (e.g. a forecaster) will only have limited data (the train and validation set). So we cannot validate our models during training on this separate test set. We can however use this test set (the full year of 2020 for BE and NL, and the full year spanning 1 July 2012 to 30 June 2013 for AUS) to do a final evaluation of the fully trained model. We split the remaining data in a train and validation set, at a proportion of 80/20.
@@
There is an important caveat to the scaling however. While it is fine to use this type of scaling on the source data, it is not as simple in the target domain if we want to apply transfer learning. The model trained on the source domain data has been scaled on 12 years of data. We can therefore reasonably assume that the out-of-sample data will not, on average, be much different from the in-sample data. However, when we naively apply a min-max scaler to the target domain data, we get data that is, at times, incompatible with the models created in the source domain. For instance, consider the case where we only have access to target data from October to December. In the Northern hemisphere, these months have the lowest PV output. So, while in reality these lagged feature values should scale to, for instance, 0.6, we scale them to 1, negatively impacting the source model performance. We tackle this issue by normalizing using the peak-power (the nominal power on the DC side in kWp) times a factor of one minus 14%, the typical percentage used by PVGIS to take into account system losses. This peak-power is known in advance for all the systems, avoiding any data leakage while still keeping the data in the correct scale for using the transfer models. Thus, the scaling is still applied as shown in equation (6), with the addition that $x_{\text{max}}$ for the target data is now:
@@ 4 Results
In this section, we present results of applying transfer learning to residential PV systems in the Netherlands and Australia, as well as the commercial PV systems in Belgium. Next, using the Dutch data, we take a closer look at the impact of time of the year on model performance. To do this, we change the training period to assess the seasonal impact. Finally, we also assess the impact of data quantity in the source domain, decreasing the amount of data available from 15 years to 3 years, and the impact of data quality by increasing the (real) distance between the source and target domain by using source data from the South of France instead of the Netherlands.
4.1 Forecasting residential PV systems
@@
The results for 40 residential PV systems in the Netherlands are summarised in figure 3. On the x-axis we plot the performance based on the months of data available in the target domain. ”0m” is the case with no target data (cf. 1. Likewise, ”1m” means that one month of target domain data used for training/finetuning. The error over time, of the baseline model (a seasonal naive model taking the values of the 24 hours of the previous day to be the new prediction), is a straight line, which does not improve over time, as it is independent of the amount of data available. On the y-axis we plot the RMSE for each of these cases. The thick lines are the overall median [^1] The median values are shown instead of mean values because not all datasets have the same observation period, which greatly impacts the mean estimate. values and the shaded areas represent the range of the individual cases. In other words, the shaded area represents the results of all 40 PV systems, from lowest to highest RMSE.
Figure 3: RMSE given months of target training data (NL)
@@
From this plot, we can make four observations. First, the transfer models strictly outperform the baseline models. Second, the target model only starts to outperform the simple baseline once a sufficient amount of data (approx. 4 to 6 months of data depending on the test site) is available to train the model. This is very likely driven by the fact that this time period is sufficient for the model to have seen months with (near) optimal PV output. Third, we can see that the asymptotic performance improvement is fairly limited once enough data becomes available. The transfer model performance only improves by about 3.3% over the months, from 0.115 RMSE to 0.111. Fourth, the transfer model almost always outperforms the target model, and continues to do so even after most of the yearly seasonality is accounted for. Figure 4 shows us, in more detail, that the transfer model outperforms the target model in 418 out of 421 scenarios. This scatter plot also shows that most of the value from transfer learning lies in the first few months when insufficient target data is available, although for some systems performance improvements can be seen even after 10 months of target data are available.
Figure 4: Delta between RMSEs of transfer and target models (more negative is in favor of the transfer model)
@@
Figure 5 plots the corresponding results for Australia. Given the larger size of the dataset (299 samples), the shaded area shows the 0.05 to 0.95 percentiles for each month to filter out some of the noise in the outliers. In a similar fashion to the Dutch results, most of the value lies in the initial months. As before, it takes over 7 months for the target model to consistently outperform the naive model here. It is worth noting that our target data to train on now ends in June, going backwards for increasing period size. Given that this is the winter period in Australia, the seasonality should not play a major role in the comparison of these two locations. The learning performance for the transfer model is again small. The median RMSE is 0.135 for the model with no fine-tuning (0m), improving by about 7.4% to 0.125 after 12 months. The target models outperform the transfer models in 8.6% of cases, mostly in the final four months.
Figure 5: RMSE given months of target training data (AUS)
4.2 Forecasting commercial PV systems
@@
Figure 6 plots the results for the two office buildings in Belgium, showing that results are in line with the results from the Netherlands. The transfer model outperforms the target model, with most of the performance difference in the first six months. The forecasting accuracy, in general, is better compared to the Netherlands. This is clear from the baseline RMSEs, which are 0.108 and 0.116 compared to the median baseline of 0.141 in the Netherlands. This is most likely due to the larger scale of the PV installation on these office buildings, as well as the diverse angles of the panels. This causes the output to be smoother compared to smaller installations with one specific angle. This in turn makes the forecasting exercise easier.
Figure 6: RMSE given months of target training data (BE)
@@ 4.3 The effect of seasonality
Next, we focus on the aspect of yearly seasonality. Given that the target models catch up to the transfer learning performance after over half a year, it is reasonable to expect that the learner requires some of the summer (or transition) months to be trained on to achieve decent performance. The training so far was done chronologically with December as the final month, i.e. for the case of ’3m’ or three months of data, the data comes from the months of October, November and December (for NL and BE). The training and validation split was also done chronologically, i.e. the validation data for this ’3m’ case consists of the last 18 days of December.
To explore this further, we shifted the terminal month in time for different models. Table 1 shows these results for NL, by focusing specifically on what happens with the RMSE of the target models relative to the daily persistence model. The columns represent the final month of the target data used for training. The earlier in the year we have the terminal month, the less data we have in total, hence the increasing amount of n.a. values. We can see that there is a clear trend: the earlier in the year the final month is, the better. This most likely has to do with the presence of the best performing day inside the training data. Out of the 40 datasets, the month with the highest PV-output hour was in April 21 times and in May eleven times. As soon as this data is available to the model, the performance drastically increases. It is also the case that the main forecasting errors come from underestimating the peaks, which further explains the need for this data to be available during the training phase.
@@
4.4 The effect of source data quantity and quality
As a final step to test the generalisation of the model, we focused on the role of the source domain and the quality of the pre-trained model. The experimental setup proposed in this paper is unique in this aspect, as pre-training on prior observed data from other similar sites does not necessarily provide this flexibility. More specifically, we assess the impact of increasing the distance between the source and target domain by using source data from the south of France. Additionally, we also look at the impact of limiting the amount of data in the source domain to three years instead of fifteen. The key results can be seen in figure 7. With regards to the data quantity, we can see that the impact is very limited (only 0.002 in scaled RMSE). If we look at the impact of changing the source location (blue line), we see a more obvious effect. The initial performance even degrades to a worse performance than the baseline. Finetuning is clearly needed in this case, as the model has to adapt to the target location, and the learning performance is more substantial here. However, the performance appears to remain worse than a model trained on the ”correct” location with a difference in the scaled RMSE of 0.012 after almost an entire year.
These results further demonstrate that i) data quantity is only important up until the point where seasonality is fully accounted for and ii) the distance between the source and target domain plays a significant role in transfer performance, even though we still see initial improvements in mis-specified source models compared to target models.
Figure 7: The RMSE over time given a limited amount of data (orange) or data from a different location (blue)

### v2

@@ 2.4 Model evaluation
Performance is measured by way of the root mean squared error (RMSE), in accordance with (Yang et al., 2020), since it penalizes large errors more heavily. This is important to correctly predict the timing and size of the daily peak in output. Another reason to use the RMSE is its ubiquity in solar forecasting literature, which makes it easier to compare across studies. In all cases, a full year of target data is kept aside for final evaluation. 80% of the remaining data is used for training and 20% is used to validate the model performance and pick the best performing iteration of the model.
@@
Figure 2: The workflow of the experiments. In the transfer learning case we train in the (synthetic) source domain before finetuning in the target domain. In direct (i.e. target) learning, we train directly on the target domain. The benchmark is a persistence model and does not require any training. The periods on the left are specifically those of the NL case, with target data starting in 2020 and ending in August 2021.
@@ 3 Data
Two types of data are required for this research: (1) the target domain data which has been observed in physical sites of interest but is often limited in quantity (and perhaps quality as well), and (2) the source domain data, which is available in ample amounts, but has actually been synthetically generated from a (similar) setup. First, we briefly discuss the target domain datasets, i.e. PV installations from the Netherlands, Australia, and Belgium, with only around 1.5 to 2 years of data per site. Second, we discuss the synthetic source domain used for transfer learning. Finally, we discuss the potential to add weather input features, as well as any necessary pre-processing steps.
@@ 3.1.1 Residential buildings in the Netherlands
For the Netherlands we have data from 40 individual household PV systems spread across the country. The data comes at a temporal resolution of 15 minutes (which we resample to an hourly resolution in accordance with the source domain). The Dutch dataset has all the required relevant site-specific meta-data available to further fine-tune our synthetic data, including the tilt, the azimuth, the peak power, and the approximate location of the PV systems (zip code and city). From these location parameters we derive an approximate latitude and longitude. The peak power of the systems varies between 1.2 kWp and 3.1 kWp, with most systems around 2.5 kWp.
For all 40 locations, the full year of 2020 is available as well as 2021 up until August. We also have a part of 2019 for several locations, but unfortunately our weather forecast data is limited to 2020 and 2021, so we cannot go back further for target locations if we want to compare univariate and multivariate performance. As mentioned, a full year of data is kept aside for evaluation of the final models. Which will thus consist of a part of 2021 and a part of 2020. This leaves little data for training and validation in the target domain, but this fits well with the core premise of the research that there is often only a (very) limited amount of data available in the target domain before forecasts need to be made.
@@ 3.1.2 Residential buildings in Australia
We also utilise solar data from 300 Australian homes with rooftop solar systems spanning the period of 1 July 2011 to 30 June 2013. The solar systems have a mean size of 1.68 kWp and are all located in New South Wales (specifically in zip codes between 20 and 23). No meta-data is available for these locations other than the size of the system and the zip code. For this dataset we will only do univariate forecasting, as no recent data is available.
@@ 3.1.3 Commercial buildings in Belgium
Finally, we also consider PV production data from two commercial buildings in Genk, Belgium. Training data is available for the full year of 2019 for one location, and for nine months in the other location (the entire year of 2020 is available for testing, which will again be univariate). In this case, the exact location of the systems is well-known, but the tilt and azimuth of these systems are not well-specified, because individual panels at the same site are placed in a variety of configurations. The systems have a peak power of 275 kWp and 125 kWp.
@@ 3.2 Source domain
For this research we use data generated by PVGIS from 2005 to 2019, mimicking the parameters (latitude, longitude, tilt, azimuth, peak power) of each of the previously discussed solar PV systems (in case these parameters are available) in order to maximize the similarity between source and target domain. So, for example, an actual PV system in EnergyVille, Genk, Belgium with latitude 50.99 and longitude 5.54 a tilt of 33°, point directly south with a peak power of 2.5 kWp would get these exact parameters in the PVGIS tool to generate data from 2005 until 2019.
This data is subsequently used to train the source models we use for transfer learning. These models are then fine-tuned using the sparse observational target data. Even though PVGIS provides data for 2020, we do not use this year while creating source models as this can potentially cause data leakage. Data leakage refers to including knowledge about the future (or data which would realistically not be available) in a model. In this case, data leakage would occur if we were to use the year 2020 in the creation of our model and, simultaneously, evaluate the model on 2020. Similarly, for the Australian PV systems, for which we have data from 2011 to 2012, we limit the source data until 2010.
3.3 Weather data
When constructing a model for predicting PV energy generation, weather data can be instrumental for accurate forecasting performance. Variables such as humidity, temperature, cloud cover, and irradiance are highly correlated with the PV power output. Ideally, one would use future weather covariates as these are best aligned with the PV energy generation. In other words, when forecasting the next 24 hours of PV, we would want to include a forecast of any of the aforementioned weather variables.
The main issue however, is access to historical weather forecasts. While historical weather data is readily available, historical forecasts are more difficult to find. They also introduce an additional dimension to the forecasting problem: the reference point at which the forecast is made. These two issues make it complex to include forecast weather data in deep learning models.
To mitigate this issue, we treat the source and target domain differently. For training the source model, we utilize readily available observed historical weather data. For the target models, we only utilize actual forecast weather data. While this introduces a mis-alignment between source and target models, it does not cause any data leakage issues and also avoids placing large and potentially expensive forecast availability requirements on SolNet users, or anyone looking to use transfer-learning for their solar power forecasts.
In this study, we retrieve historical weather data from Open-Meteo via its API. The variables used in this study are relative humidity, direct irradiance, and diffuse irradiance. All that is required is to specify a date-range (which is 2005 until 2019 for our source model), and a latitude and longitude. We did find some open-source historical weather forecasts, but only from 2020 onwards, and only for Europe. The data we use comes from DWD ICON. [^1] More details can be found on the website of the DWD: https://www.dwd.de/EN/Home/home_node.html They maintain both a global and a European Numerical Weather Prediction (NWP) model and make the data publicly available. However, the forecasts are only kept available for a 24 hour period, but Open Climate Fix maintains a repository on Hugging Face for the EU data going back until 2020 (Open Climate Fix, 2023). This data is used to fine-tune the transfer model on our target domain and do inference in the target domain. In doing so we can show that using historical weather to train our source models is sufficient to gain a good increase in performance, allowing people to use deep learned models including weather data, which is still easily accessible (through PVGIS and Open-Meteo).
@@ 3.4 Data pre-processing
The source data is split into a train and test set, while the target data is split into a train set, a validation set, and a separate test set. We do this three-way split to mimic a realistic scenario where a user of the system (e.g. a forecaster) will only have limited data (the train and validation set). So we cannot validate our models during training on this separate test set. We can however use this test set of one year to do a final evaluation of the fully trained model. We split the remaining data in a train and validation set, at a proportion of 80/20.
@@
There is an important caveat to the scaling however. While it is fine to use this type of scaling on the source data, it is not as simple in the target domain if we want to apply transfer learning. The model trained on the source domain data has been scaled on 12 years of data. We can therefore reasonably assume that the out-of-sample data will not, on average, be much different from the in-sample data. However, when we naively apply a min-max scaler to the (at times very limited) target domain data, we get data that is, at times, incompatible with the models created in the source domain. For instance, consider the case where we have PV data of a winter month in the Netherlands. When applying a min-max scaler to this unrepresentative dataset, we would scale the values between 0 and 1. But values from summer months would, in that case, go up to 2 or 3, while these values are scaled to 1 in the source model. We tackle this issue by normalizing the data using domain knowledge, i.e. the peak-power (the nominal power on the DC side in kWp) times a factor of one minus 14% (the typical percentage used by PVGIS to take into account system losses). This peak-power is known in advance for all the systems, avoiding any data leakage while still keeping the data scaled accurately for use in transfer models. Thus, the scaling is still applied as shown in equation (6), with the addition that $x_{\text{max}}$ for the target data is now:
@@ 4 Results
In this section, we present results of applying transfer learning to residential PV systems in the Netherlands and Australia, as well as the commercial PV systems in Belgium using univariate data. Next, using the Dutch data, we take a closer look at the impact of including weather covariates. Then we look at the importance of the time of the year when we start gathering target data on model performance. To do this, we change the training period to assess the seasonal impact. Finally, we also assess the impact of data mis-specification in the source domain, by increasing the (real) distance between the source and target domain.
4.1 Univariate forecasting of residential PV systems
@@
The results for univariate forecasting of 40 residential PV systems in the Netherlands are summarised in figure 3(a) (i.e. without incorporating weather observations or forecasts in either the source or target domain). On the x-axis we plot the performance based on the months of data available in the target domain. ”0m” is the case with no target data (cf. 1. Likewise, ”1m” means that one month of target domain data is used for training/finetuning. On the y-axis we plot the scaled RMSE for each of these cases (i.e. the RMSE on the data which has been scaled using the min-max scaler). The thick lines are the overall median [^2] The median values are shown instead of mean values because not all datasets have the same observation period, which greatly impacts the mean estimate. values and the shaded areas represent the range of the individual cases. In other words, the shaded area represents the results of all 40 PV systems, from lowest to highest RMSE. The error of the baseline model (a seasonal naive model taking the values of the 24 hours of the previous day to be the new prediction), is a straight line, which does not improve over time, as it is independent of the amount of data available.
(a) NL scaled RMSE
(b) AUS scaled RMSE
(c) BE scaled RMSE
Figure 3: Forecast error as a function of target data availability and use of transfer learning for univariate models. Results indicate transfer learning models consistently outperform both naive baselines and target models (i.e. trained without transfer)
@@
From this plot, we can make four observations. First, the transfer models strictly outperform the baseline models. Second, the target model only starts to outperform the simple baseline once a sufficient amount of data (approx. 4 to 6 months of data depending on the test site) is available to train the model. This is very likely driven by the fact that this time period is sufficient for the model to have seen months with different weather conditions allowing it to generalize. Third, we can see that the asymptotic performance improvement is almost non-existent, with the RMSE staying at roughly 0.115. Fourth, the transfer model almost always outperforms the target model, until more than a half year of data is available.
@@
Figure 3(b) plots the corresponding results for Australia. Given the larger size of the dataset (299 samples), the shaded area shows the 0.05 to 0.95 percentiles for each month to filter out some of the noise in the outliers. In a similar fashion to the Dutch results, most of the value lies in the initial months. As before, it takes over 7 months for the target model to consistently outperform the naive model here. The learning improvement performance for the transfer model is again small. The median RMSE is 0.135 for the model with no fine-tuning (0m), improving by about 7.4% to 0.125 after 12 months. The target models outperform the transfer models in 8.6% of cases, mostly in the final four months.
4.2 Univariate forecasting of commercial PV systems
@@
Figure 3(c) plots the results for the two commercial buildings in Belgium, showing that results are in line with the residential results. As before, the transfer model outperforms the target model, with most of the performance difference in the first six months. The transfer forecasting accuracy is slightly better compared to the Dutch case. This is clear from the baseline RMSEs, which are 0.108 and 0.116 compared to the median baseline of 0.141 in the Netherlands. This is most likely due to the larger scale of the PV installation, as well as the diverse angles of the panels. This causes the output to be smoother compared to smaller installations with one specific angle, making the forecasting exercise easier.
A major difference between BE and NL, however, is that the final month of training lies in the winter for BE and in summer for NL. This would explain why it takes so much longer for a univariate target model in BE to perform decently compared to NL, because it does not have any data from summer months to fine-tune on for the first few months. (5m is August in this case, 6m is July) We will come back to this issue in section 4.4.
4.3 The effect of weather covariates
Figure 4 shows the results for the Dutch buildings, but now with the addition of three weather variables: relative humidity at 2 meters, direct solar radiation, and diffuse solar radiation. From the plot, we can come to several conclusions. First, the performance increases significantly when including weather variables, the median transfer error decreases from c. 0.117 to 0.085, a performance increase of 27.4%. Second, the performance increase is even more pronounced with target models. When no covariates were included, 7 months of data were needed by the target LSTM to outperform even the naive model on average, but with weather data it only takes 2 months. On average, the transfer model still performs better for the first half year, but the difference is smaller compared to the univariate case. A final observation is that the performance is far more consistent across different sites now, i.e. the confidence band of the target and transfer models is a lot narrower now.
Figure 4: Forecast error as a function of target data availability and use of transfer learning for multivariate models using weather data. The target model is trained on weather forecast data, the transfer model is trained on actuals weather data and fine-tuned on weather forecast data. Results are consistent with the univariate models, but with a strong increase in overall performance.
@@ 4.4 The effect of seasonality
Next, we focus on the aspect of yearly seasonality. Given that the target models catch up to the transfer learning performance after over half a year, it is reasonable to expect that the learner requires some of the summer (or transition) months to be trained on to achieve decent performance. The training so far was done chronologically with July as the final month for NL and AUS, and December as the final month for BE, i.e. for the case of ’3m’ or three months of data, the data comes from the months of October, November and December for BE. The training and validation split was also done chronologically, i.e. the validation data for this ’3m’ case consists of the last 18 days of December.
To explore this further, we shifted the terminal month in time for different univariate, models. Table 1 shows these results for NL, by focusing specifically on what happens with the RMSE of the target models relative to the daily persistence model. The columns represent the final month of the target data used for training. The earlier in the year we have the terminal month, the less data we have in total, hence the increasing amount of n.a. values. We can see that there is a clear trend: the earlier in the year the final month is, the better. This most likely has to do with the presence of the best performing day inside the training data. Out of the 40 datasets, the month with the highest PV-output hour was in April 21 times and in May eleven times. As soon as this data is available to the model, the performance drastically increases. It is also the case that the main forecasting errors come from underestimating the peaks, which further explains the need for this data to be available during the training phase.
@@
4.5 The effect of source data mis-specification
As a final step to test the generalisation of the model, we focused on the role of the source domain and the quality of the pre-trained model. The experimental setup proposed in this paper is unique in this aspect, as pre-training on prior observed data from other similar sites does not necessarily provide this flexibility. More specifically, we assess the impact of increasing the distance between the source and target domain by using source data increasingly to the south-east of the original sites (at an angle of 125° to avoid ending up in the Adriatic Sea after 800km). Figure 5 plots these results for a subset of ten locations of our NL dataset. Again, the lines show the median values, while the shaded area shows the confidence interval.
Two key findings become clear from these figures. On the one hand, we can see that mis-specification of the target location comes with increasingly erroneous forecasts. On the other hand, the use of external covariates plays a key role in the impact of these mis-specifications. When weather covariates are included in the model (Figure 5(b)), the results worsen drastically with increased distance from the target location. Moving 100km to the south-east increases the error by 15% (from 0.080 to 0.093) over the correct location, 35% at 200km (from 0.080 to 0.108) and at 400km the error is already at a similar level to a naive model. When no weather variables are included (Figure 5(a)), this effect is far less pronounced. The performance definitely lags behind on a model with weather covariates, but the model is much more robust towards mis-specifications.
The median target RMSE (the blue line) also makes clear how important correct specification of the target location is when weather data is included. When the location is mis-specified by c. 200km we can see that the transfer model has already lost all its relative value after two months. When the mis-specification is as small as 100km, the value is lost after 5 months. However, when no weather data is included, a mis-specification of the source domain has almost no effect on target outperformance, even after more than 6 months.
(a) Without weather covariates
(b) With weather covariates
Figure 5: Transfer learning results under increasing distance from the target for a subset of ten NL locations. The performance is compared to the median baseline of these ten locations, as well as the median ’correct’ target model. I.e. a target model that is made with no mis-specification.